OpenStudy (anonymous):

Choose the correct simplification of the expression (a^2overb^5) cubed

OpenStudy (skullpatrol):

$$(\frac{a}{b})^m=\frac{a^m}{b^m}$$

OpenStudy (anonymous):

so a^5 overb^8

OpenStudy (skullpatrol):

$$(a^m)^n=a^{mn}$$

OpenStudy (anonymous):

i Cant see that

OpenStudy (skullpatrol):